OpenStudy (amistre64):

how about; why is: Dx(e^x) = e^x : considered a seperate rule from general exponential derivatives?

OpenStudy (anonymous):

U should be careful, u will be turned into an analyst before u know it....:-)

OpenStudy (amistre64):

::: looks to see if its a full moon :::

OpenStudy (amistre64):

\[[B^x]'=B^x\ ln(B)\ x'\] \[[e^x]'=e^x\ ln(e)\ x'\] right?

OpenStudy (zarkon):

yes

OpenStudy (amistre64):

why do we need 2 rules for this one then?

myininaya (myininaya):

lol

OpenStudy (zarkon):

you don't

myininaya (myininaya):

B can equal e

OpenStudy (amistre64):

my teacher said e was a special case ...

myininaya (myininaya):

amistre what are you doing to yourself?

OpenStudy (anonymous):

Consult generalized power rules....

OpenStudy (amistre64):

im staving off going comatose from the onslaught of algbra

OpenStudy (zarkon):

only because ln(e)=1

myininaya (myininaya):

why does 1=0 again zarkon?

OpenStudy (anonymous):

are you supposed to prove that the derivative of \[e^x\] is \[e^x\]?

OpenStudy (zarkon):

lol...don't thin i want to go there ;)

myininaya (myininaya):

lol

OpenStudy (anonymous):

i think this depends largely on your definition of e

myininaya (myininaya):

i thought amistre has taken calculus

OpenStudy (anonymous):

Rule 1 - I f something doesn't work properly, redefine it. Rule 2 - I f it still doesn't work, generalize it.

OpenStudy (amistre64):

i taught myself calculus; going thru the courses now tho

myininaya (myininaya):

wow amistre you are so smart

OpenStudy (anonymous):

one definition is that it is the number that works. you can see that the derivative of \[2^x=2^x\lim_{h\rightarrow 0}\frac{2^h-1}{h}\]

myininaya (myininaya):

theres no way i could have done that

OpenStudy (amistre64):

... i wasnt smart enough to know 0^0 was bad lol

OpenStudy (anonymous):

and the derivative of \[3^x=3^x\lim_{h\rightarrow 0}\frac{3^h-1}{h}\]

OpenStudy (amistre64):

why is the derivative of e^x a seperate rule from B^x, is my question in general

OpenStudy (anonymous):

e^x came from logs....

myininaya (myininaya):

there is no reason they should be separate

OpenStudy (anonymous):

one limit is less than 1, one is greater, and you can define e to be the number such that \[\lim_{h\rightarrow 0}\frac{e^h-1}{h}=1\]

OpenStudy (anonymous):

that gives it to you.

myininaya (myininaya):

think they talk about the rule d(e^x)=e^x first because it is easiest

OpenStudy (anonymous):

actually i think e came from exponential growth, not from logs but i might be mistaken. pretty sure though. euler looking at continuous growth

myininaya (myininaya):

i want a potato

OpenStudy (anonymous):

Logs, very interesting history, u would like it....

OpenStudy (anonymous):

starting with napier?

OpenStudy (anonymous):

Yup...

OpenStudy (anonymous):

potato mmmm too many carbs for me

OpenStudy (anonymous):

Working it in as an inverse came later...

OpenStudy (amistre64):

i tend to start earlier; with thag and goor :)

OpenStudy (anonymous):

used as calculator as i remember but not really logs as we think of them

OpenStudy (amistre64):

caveman maths

OpenStudy (anonymous):



OpenStudy (anonymous):

e was there, just didn't calculate it as e.

OpenStudy (anonymous):

napier? i need a book

OpenStudy (anonymous):

@estudier did you prove 1/4 is in the cantor set?

OpenStudy (amistre64):

napier and...bernolli ?

OpenStudy (anonymous):

must have been some bernoulli. they did everything.

OpenStudy (anonymous):

1/4 is in the ternary cantor set (obviously)